# Scalable Language Models with Posterior Inference of Latent Thought Vectors

### v1

@@ Abstract
We propose a novel family of language models, Latent-Thought Language Models (LTMs), which incorporate explicit latent thought vectors that follow an explicit prior model in latent space. These latent thought vectors guide the autoregressive generation of ground tokens through a Transformer decoder. Training employs a dual-rate optimization process within the classical variational Bayes framework: fast learning of local variational parameters for the posterior distribution of latent vectors, and slow learning of global decoder parameters. Empirical studies reveal that LTMs possess additional scaling dimensions beyond traditional LLMs, yielding a structured design space. Higher sample efficiency can be achieved by increasing training compute per token, with further gains possible by trading model size for more inference steps. Designed based on these scaling properties, LTMs demonstrate superior sample and parameter efficiency compared to conventional autoregressive models and discrete diffusion models. They significantly outperform these counterparts in validation perplexity and zero-shot language modeling. Additionally, LTMs exhibit emergent few-shot in-context reasoning capabilities that scale with model and latent size, and achieve competitive performance in conditional and unconditional text generation.
@@ 1 Introduction
Figure 1: Design space of Latent-Thought Language Models (LTMs), Autoregressive Models (ARMs) and Masked Diffusion Models (MDMs). LTMs introduce additional scaling dimensions.
@@
The remarkable success of LLMs is underpinned by well-established scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), which predict performance improvements with increased model size, data size, and training compute. The induced equations reveal that larger models achieve significantly higher sample efficiency (evaluated by the number of training tokens for achieving certain performance), making it computationally optimal to train very large models and stop before convergence. However, recent observations of diminishing returns from scaling in gigantic models have led researchers to explore alternative scaling dimensions (Snell et al., 2024). We are thus motivated to explore a novel family of language models that expand these possibilities.
@@
We propose Latent-Thought Language Models (LTMs), which incorporate explicit latent thought vectors that follow explicit prior model in the latent space. These latent vectors control an autoregressive Transformer decoder’s (Vaswani et al., 2017) generation of each token throughout the sequence, effectively creating an abstract representation of the entire sequence. LTMs are trained within the classical variational Bayes framework (Jordan et al., 1999; Blei et al., 2017; Murphy, 2012), with a dual-rate optimization process: fast learning of local variational parameters for the posterior distribution of latent thought vectors, coupled with slow learning of global decoder parameters. This approach enables rapid adaptation to specific inputs while gradually accumulating general linguistic knowledge.
@@
LTMs’ architecture and learning scheme are inspired by established cognitive models. Within the framework of the declarative-procedural model (Ullman, 2004), the latent thought vectors and local variational parameters parallel the declarative or episodic memory, while the global decoder parameters correspond to procedural memory. The dual-rate learning scheme reflects the interplay between fast episodic learning and slow schematic learning in human cognition (Kumaran et al., 2016). Furthermore, in the context of language of thought hypothesis (Fodor, 1975), the latent thought vectors can be interpreted as “words” of a language of thought.
@@
Beyond their cognitive science foundations, LTMs introduce two novel dimensions for investigating scaling behaviors: inference steps and latent size. To empirically study the scaling behaviors of LTMs, we conducted extensive experiments at GPT-2 scale (Radford et al., 2019) using the OpenWebText dataset (Gokaslan & Cohen, 2019). The perplexity of LTMs scales with data size, model size, inference steps, and latent size. These factors form a hierarchical decision tree as shown in Figure 1. While traditional LLMs primarily trade off between data size and model size, LTMs introduce a higher-level trade-off among data size, training FLOPs per token (trFLOPs/tok), and latent size. At a fixed latent size, increased trFLOPs/tok yields higher sample efficiency. Increasing latent size reduces sample efficiency but enhances reasoning capabilities with minimal impact on trFLOPs/tok. Once trFLOPs/tok is determined, we can optimize the trade-off between inference steps and model size: more inference steps improves sample and compute efficiency until a bottleneck. These relationships provide preliminary guidance for sample-efficient and compute-optimal training of LTMs.
@@
In comparison with traditional autoregressive models (Radford et al., 2019) and more recent diffusion-based approaches (Lou et al., 2024; Shi et al., 2024; Sahoo et al., 2024), LTMs demonstrate superior efficiency in data and parameters, and excel in several key NLP tasks:
@@
  • Arithmetic reasoning: LTMs demonstrate emergent few-shot in-context learning at scales that are significantly smaller than GPTs. This is significant even in our smallest model, LTM-Small. This capability scales further with increased model and latent size.
@@
Contributions. Language models with explicit latent thought vectors that follow a prior model in latent space are much under-explored in recent years. Compared to ground tokens, the latent thought vectors provide a highly compact, abstract and structured representation in a lifted latent space. Our paper constitutes a systematic exploration of this model class with the following contributions:
@@
    Comprehensive analysis of scaling properties, especially along the dimensions of inference steps and latent size.
@@
Our work lays the groundwork for future development of this model class.
@@ 2 Method
2.1 Latent-Thought Language Models (LTMs)
@@
We formulate our framework as a structured probabilistic model that captures the relationship between latent thought vectors and observed sequences as shown in Figure 2.
Layered Thought Vectors. We assume ${\mathbf{z}}={({\mathbf{z}}_{1},...,{\mathbf{z}}_{L})}$, where ${\mathbf{z}}_{l}$ consists of thought vectors cross-attending to layer $l$ of the Transformer decoder. While we explored an alternative design using a single set of thought vectors attending to all layers simultaneously, empirical evidence strongly favors the layered approach. The layered structure, where distinct sets of thought vectors attend to different layers, appears to capture multiple levels of abstraction more effectively.
Prior Model. For the prior model $p({\mathbf{z}})$, we assume an isotropic Gaussian prior over the latent thought vectors ${\mathbf{z}}={({\mathbf{z}}_{1},...,{\mathbf{z}}_{L})}\sim\mathcal{N}(\mathbf{0},\bf{I})$. We have also experimented with more complex alternatives, such as a neural transport model ${\mathbf{z}}=f_{\alpha}({\mathbf{z}}_{0})$, with ${\mathbf{z}}_{0}\sim\mathcal{N}(\mathbf{0},\bf{I})$, and $f_{\alpha}$ being a Transformer encoder. However, empirically we find the simple Gaussian prior works well so we adopt it as the default.
@@
Short Context Window. We are particularly interested in models with a short context window of size $k$: $p_{\beta}({\mathbf{x}}|{\mathbf{z}})=\prod_{n=1}^{N}p_{\beta}(x^{(n)}|{\mathbf{z}},{\mathbf{x}}^{(n-k:n-1)}),$ where ${\mathbf{x}}^{(n-k:n-1)}$ denotes the $k$ previous elements. This short context forces ${\mathbf{z}}$ to serve as a global information carrier, integrating information across temporal segments that would otherwise be disconnected due to the short context window. $k=256$ in our experiments.
@@ 2.2 Learning and Posterior Inference
It is crucial to emphasize that $(\bm{\mu},\bm{\sigma}^{2})$ are local parameters, specific to each training or testing sequence ${\mathbf{x}}$. This is in contrast to the parameters in the decoder generator, which are shared by all the training sequences and thus are global parameters. As detailed in Algorithm 1, we employ a dual-rate learning algorithm: fast inference of local parameters using a gradient descent algorithm, AdamW (Kingma & Ba, 2014; Loshchilov & Hutter, 2019), with high learning rates (e.g., 0.3) and few steps (e.g., 16), alternating with slow updates of global decoder parameters (e.g., learning rate 0.0004). This enables rapid per-instance adaptation while gradually building general linguistic knowledge.
In our work, we use finite number of steps (e.g., $T_{\text{fast}}=16$) for posterior inference or fast learning. Such a finite-step inference compute is usually affordable on modern GPUs, especially for a relatively small decoder model with short context window. The joint optimization of ELBO (5) has the effect that the learned decoder model will bias itself from maximum likelihood to accommodate this finite-step inference to make the resulting variational posterior close to the true posterior. It is well known (Murphy, 2012) that variational inference has mode chasing behavior, so making $T$ too large may cause the variational posterior to overfit the training example in the sense that it becomes overly confident, which may lead to underfitting of the decoder model. In our experiments, we shall study the scaling behavior along the dimension of $T_{\text{fast}}$.
@@
Variational Autoencoder with Amortized Inference. As another baseline, the VAE approach (Kingma & Welling, 2013) introduces an inference model $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ with global parameters $\phi$ to amortize the iterative inference computation in classical variational learning. In our experiments on VAE, we observe severe posterior collapse (Lucas et al., 2019; Pang et al., 2021), even with careful annealing on the KL-divergence term in ELBO (5). The reason is that the number of global parameters in the inference model is fixed, whereas the number of local parameters in the classical variational Bayes inference is proportional to the number of training examples. As a result, it is much easier for the classical variational Bayes method to minimize the ELBO with such a large number of local parameters, whereas it is very tempting for the inference model to go the easy route to minimize only the KL term in ELBO (5). A simple fix is to infer the local parameters in the traditional variational Bayes framework, and then distill the inferred local parameters to the inference model, which serves as a memorization. We can integrate inference model into classical variational Bayes framework, where the inference model memorizes the classical variational inference results while jump-starting the classical variational inference. We will leave this cooperative scheme to future work.
@@
Comparisons. We adopt classical variational Bayes, leaving Langevin-based learning and VAE as ablation baselines. Compared to Langevin sampling, it provides more efficient optimization. Compared to VAE, it avoids learning a large inference model and mitigates posterior collapse by avoiding the initial mismatch between the inference model and the true posterior. More importantly, the classical variational method allows us to explore gradient descent for inference, connecting our approach to fast-slow learning and test-time computation paradigms (Ba et al., 2016; Krause et al., 2018).
Algorithm 1 Dual-rate learning of LTM
@@
5:        // fast learning
@@
10:           Update $\bm{\mu}_{i},\bm{\sigma}^{2}_{i}$ using AdamW with $\eta_{\text{fast}}$.
@@ 2.3 Conditional and Unconditional Generation
We sample the posterior distribution $p({\mathbf{z}}|{\mathbf{x}})\propto p({\mathbf{z}})p_{\beta}({\mathbf{x}}|{\mathbf{z}})$ using classical variational inference, following the same mechanism as the fast learning of $q({\mathbf{z}}|{\mathbf{x}})$ in Equation 5 during training. The actual sampling distribution becomes:
@@
Zelikman et al. (2022); Hu et al. (2023); Hoffman et al. (2024) also sample posterior latent (chain-of-)thoughts for conditional generation from $p({\mathbf{y}}|{\mathbf{x}})$, but their approaches differ fundamentally from LTMs since they work on post-training of traditional autoregressive models on finetuning sets, while LTMs’ posterior inference is naturally optimized during pre-training. Sampling from $p_{\beta}({\mathbf{y}}|{\mathbf{x}},{\mathbf{z}})$ follows standard autoregressive sampling techniques (Freitag & Al-Onaizan, 2017; Holtzman et al., 2019). For unconditional generation, we sample from:
@@
An alternative sampling scheme involves incorporating each newly generated token into the prefix and then updating ${\mathbf{z}}$ through variational inference. We leave exploration of this more computationally intensive approach to future work.
@@
2.4 Inference Compute
Compared to language models operating in the token space (e.g., ARMs and MDMs), LTMs introduce a distinct computational cost in the form of inference compute — a requirement stemming from the fast learning of latent thought vectors. This inference compute occurs in both model training and testing. Let’s start from analyzing it within the context of total training compute.
For one nested iteration of LTM’s dual-rate learning with $T_{\text{fast}}$ inference steps on an input sequence of $N$ tokens (vocabulary size $V$), we consider a model with $L$ attention layers, $N_{{\mathbf{z}}}$ latent thought vectors per layer, and hidden dimension $H$. The forward pass computational complexity is approximately $\mathcal{O}(L(N^{2}H+NN_{{\mathbf{z}}}H+NH^{2})+NVH)$, comprising $\mathcal{O}(N^{2}H)$ for self-attention, $\mathcal{O}(NN_{{\mathbf{z}}}H)$ for cross-attention with latent vectors, $\mathcal{O}(NH^{2})$ for feed-forward layers, and $\mathcal{O}(NVH)$ for embedding layers. The backward pass doubles this cost due to gradient computation and activation storage (Chowdhery et al., 2023). With $T_{\text{fast}}$ backward passes in fast learning, and $1$ additional backward pass in slow learning, the training compute per token (trFLOPs/tok) is $\mathcal{O}((T_{\text{fast}}+1)L(NH+N_{{\mathbf{z}}}H+H^{2})+(T_{\text{fast}}+1)VH)$. Thus, while both LTMs and ARMs involve gradient back-propagation for training, LTMs distribute compute differently: they trade ARMs’ compute in slow learning of global parameters for fast learning of local parameters.
Among scaling factors $N_{{\mathbf{z}}}$, $L$, and $T_{\text{fast}}$, $N_{{\mathbf{z}}}$ minimally impacts trFLOPs/tok since latent vectors are more compressed than sequence length ($N_{{\mathbf{z}}}\ll N$). $L$’s impact exceeds embedding layers only when $L>V/N$, as illustrated in Figure 3. $T_{\text{fast}}$ most significantly influences trFLOPs/tok. Notably, when $T_{\text{fast}}\gg 1$, the compute for fast learning subsumes slow learning, and the trFLOPs/tok of $\mathcal{O}(T_{\text{fast}}L(NH+N_{{\mathbf{z}}}H+H^{2})+T_{\text{fast}}VH)$ represents both the training compute (with slow learnig) and the inference compute (without slow learning).
During testing, $N$ varies by task: it represents the text length for latent thought inference in likelihood estimation and generation tasks. As detailed in Section 2.3, generation tasks’ inference compute can further vary by sampling scheme. For our adopted sampling scheme, trFLOPs/tok provides a worst-case estimate of inference compute across all tasks.

### v2

# Latent Thought Models with Variational Bayes Inference-Time Computation

@@ Abstract
We propose a novel class of language models, Latent Thought Models (LTMs), which incorporate explicit latent thought vectors that follow an explicit prior model in latent space. These latent thought vectors guide the autoregressive generation of ground tokens through a Transformer decoder. Training employs a dual-rate optimization process within the classical variational Bayes framework: fast learning of local variational parameters for the posterior distribution of latent vectors (inference-time computation), and slow learning of global decoder parameters. Empirical studies reveal that LTMs possess additional scaling dimensions beyond traditional Large Language Models (LLMs), such as the number of iterations in inference-time computation and number of latent thought vectors. Higher sample efficiency can be achieved by increasing training compute per token, with further gains possible by trading model size for more inference steps. Designed based on these scaling properties, LTMs demonstrate superior sample and parameter efficiency compared to autoregressive models and discrete diffusion models. They significantly outperform these counterparts in validation perplexity and zero-shot language modeling tasks. Additionally, LTMs exhibit emergent few-shot in-context reasoning capabilities that scale with model size, and achieve competitive performance in conditional and unconditional text generation. The project page is available at https://deqiankong.github.io/blogs/ltm.
@@ 1 Introduction
Figure 1: Analysis of model scaling behavior of validation perplexity across model size, inference steps, and the number of latent thought vectors $N_{\mathbf{z}}$. Autoregressive and diffusion baselines are plotted as dashed lines.
@@
The remarkable success of LLMs is underpinned by well-established scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), which predict performance improvements with increased model and data size. The induced equations reveal that larger models achieve significantly higher sample efficiency (evaluated by the number of training tokens for achieving certain performance), making it computationally optimal to train very large models and stop before convergence. However, as model sizes grow rapidly, data availability has emerged as a critical bottleneck for continued scaling. This limitation motivates our exploration of a novel class of language models that introduces new scaling dimensions to unlock further improvements in sample efficiency.
@@
We propose Latent Thought Models (LTMs), which incorporate explicit latent thought vectors that follow explicit prior model in the latent space. These latent vectors control an autoregressive Transformer decoder’s (Vaswani et al., 2017) generation of each token throughout the sequence, effectively creating an abstract representation of the entire sequence. LTMs are trained within the classical variational Bayes framework (Jordan et al., 1999; Blei et al., 2017; Murphy, 2012), with a dual-rate optimization process: fast learning or inference-time computation of local variational parameters for the posterior distribution of latent vectors, and slow learning of global decoder parameters. This approach enables rapid adaptation to specific inputs while gradually accumulating general linguistic knowledge.
@@
The architecture and learning scheme of LTMs draw inspiration from established cognitive models. Within the framework of the declarative-procedural model (Ullman, 2004), the latent thought vectors and local variational parameters parallel the declarative or episodic memory, while the global decoder parameters correspond to procedural memory. The dual-rate learning scheme reflects the interplay between fast episodic learning and slow schematic learning in human cognition (Kumaran et al., 2016). Moreover, under the language of thought hypothesis (Fodor, 1975), the latent thought vectors can be interpreted as “words” of an internal cognitive language.
@@
LTMs introduce novel dimensions for investigating scaling behaviors: the number of iterations in inference-time computation (inference steps), and the number of latent thought vectors (latent size). To empirically study the scaling behaviors of LTMs, we conducted extensive experiments at GPT-2 scale (Radford et al., 2019) using the OpenWebText dataset (Gokaslan & Cohen, 2019). The perplexity of LTMs scales with data size, model size, inference steps and latent size. While traditional LLMs primarily trade off between data size and model size, LTMs introduce a higher-level trade-off between data size and compute per token ( training FLOPs per token (trFLOPs/tok)). At a fixed trFLOPs/tok budget, LTMs can be optimized across multiple dimensions: inference steps, model size, and latent size. While scaling any of these dimensions improves performance, as shown in Fig. 1, increasing inference steps enhances both sample and compute efficiency, with larger latent sizes providing additional headroom for improvement (Fig. 4). These relationships provide preliminary guidance for sample-efficient and compute-optimal training of LTMs, revealing that inference-time computation represents a fundamentally new axis that complements traditional model parameter and data scaling.
@@
In comparison with traditional autoregressive models (Radford et al., 2019) and more recent diffusion-based approaches (Lou et al., 2024; Shi et al., 2024; Sahoo et al., 2024), LTMs demonstrate superior efficiency in data and parameters, and excel in several key language tasks:
@@
  • Arithmetic Reasoning: LTMs demonstrate emergent few-shot in-context learning at scales that are significantly smaller than GPTs. This is significant even in our smallest model, LTM-Small. This capability scales further with increased model size. We also find scaling the number of latent thought vectors appears to be helpful.
@@
Contributions. Language models with explicit latent thought vectors that follow a prior model in latent space are much under-explored in recent years. Compared to ground tokens, the latent thought vectors provide a highly compact, abstract and structured representation in a lifted latent space. This paper constitutes a systematic exploration of this model class with the following contributions:
@@
    Comprehensive analysis of scaling properties, especially along the dimensions of inference steps and model size.
@@ 2 Method
2.1 Latent Thought Models (LTMs)
@@
We formulate our framework as a structured probabilistic model that captures the relationship between latent thought vectors and observed sequences as shown in Fig. 2.
Layered Thought Vectors. We assume ${\mathbf{z}}={({\mathbf{z}}_{1},...,{\mathbf{z}}_{L})}$, where ${\mathbf{z}}_{l}$ consists of thought vectors cross-attending to layer $l$ of the Transformer decoder. $N_{\mathbf{z}}$ denotes the total number of latent vectors, except in Section 2.4 where it represents the number per layer. While we explored an alternative design using a single set of thought vectors attending to all layers simultaneously, empirical evidence strongly favors the layered approach. The layered structure, where distinct sets of thought vectors attend to different layers, appears to capture multiple levels of abstraction more effectively.
Prior Model. For the prior model $p({\mathbf{z}})$, we assume an isotropic Gaussian prior over the latent thought vectors ${\mathbf{z}}={({\mathbf{z}}_{1},...,{\mathbf{z}}_{L})}\sim\mathcal{N}(\mathbf{0},\bf{I})$. This prior model is a proper starting point due to its simplicity. It is already a structured prior model with multiple layers of latent thought vectors. We shall explore more sophisticated learnable prior model $p_{\alpha}({\mathbf{z}})$ in future work.
@@
Short Context Window. We are particularly interested in models with a short context window of size $k$: $p_{\beta}({\mathbf{x}}|{\mathbf{z}})=\prod_{n=1}^{N}p_{\beta}(x^{(n)}|{\mathbf{z}},{\mathbf{x}}^{(n-k:n-1)}),$ where ${\mathbf{x}}^{(n-k:n-1)}$ denotes the $k$ previous elements. This short context forces ${\mathbf{z}}$ to serve as a information carrier, integrating information across temporal segments that would otherwise be disconnected due to the short context window. $k=256$ in our experiments.
@@ 2.2 Learning and Posterior Inference
It is crucial to emphasize that $(\bm{\mu},\bm{\sigma}^{2})$ are local parameters, specific to each training or testing sequence ${\mathbf{x}}$. This is in contrast to the parameters in the decoder generator, which are shared by all the training sequences and thus are global parameters. As detailed in Algorithm 1, we employ a dual-rate learning algorithm: fast inference of local parameters using a gradient descent algorithm, Adam (Kingma & Ba, 2014; Loshchilov & Hutter, 2019), with high learning rates (e.g., 0.3) and few steps (e.g., 16), alternating with slow updates of global decoder parameters (e.g., learning rate 0.0004). This enables rapid per-instance adaptation while gradually building general linguistic knowledge.
In our work, we use finite number of steps (e.g., $T_{\text{fast}}=16$) for fast learning or inference-time computation for the posterior distribution of latent thought vectors. Such a finite-step inference-time computation is usually affordable on modern GPUs, especially for a relatively small decoder model with short context window. While finite-step fast learning may introduce a bias relative to maximum likelihood if local variational inference does not converge (Hoffman et al., 2013), we empirically study how scaling the number of steps influences this bias under LTMs’ architectural conditions.
@@
Variational Autoencoder with Amortized Inference. As another baseline, the VAE approach (Kingma & Welling, 2013) introduces an inference model $q_{\phi}({\mathbf{z}}|{\mathbf{x}})$ with global parameters $\phi$ to amortize the iterative inference computation in classical variational learning. In our experiments on VAE, we observe severe posterior collapse (Lucas et al., 2019; Pang et al., 2021), even with careful annealing on the KL-divergence term in ELBO (Eq. 5). Note that the inference model only has a fixed number of parameters, which are shared by all data points, while the classical variational Bayes inference has local parameters whose size is proportional to the number of training examples. As a result, the inference model is more likely than the classical variational Bayes to take the easy route and only minimize the KL term in ELBO. A simple fix is to infer the local parameters in the traditional variational Bayes framework, and then distill the inferred local parameters to the inference model.
@@
Comparisons. We adopt classical variational Bayes, leaving Langevin-based learning and VAE as ablation baselines. Compared to Langevin sampling, it provides more efficient optimization. Compared to VAE, it avoids learning a large inference model and mitigates posterior collapse by avoiding the initial mismatch between the inference model and the true posterior. More importantly, the classical variational method allows us to explore gradient descent for inference, connecting our approach to fast-slow learning and inference-time or test-time computation paradigms (Ba et al., 2016; Krause et al., 2018).
Algorithm 1 Fast-Slow Learning of LTM
@@
5:        // fast learning or Inference-time computation
@@
10:           Update $\bm{\mu}_{i},\bm{\sigma}^{2}_{i}$ using Adam with $\eta_{\text{fast}}$.
@@ 2.3 Conditional and Unconditional Generation
We sample the posterior distribution $p({\mathbf{z}}|{\mathbf{x}})\propto p({\mathbf{z}})p_{\beta}({\mathbf{x}}|{\mathbf{z}})$ using classical variational inference, following the same mechanism as the fast learning of $q({\mathbf{z}}|{\mathbf{x}})$ in Eq. 5 during training. The actual sampling distribution becomes:
@@
Zelikman et al. (2022); Hu et al. (2023); Phan et al. (2023) also sample posterior latent (chain-of-)thoughts for conditional generation from $p({\mathbf{y}}|{\mathbf{x}})$, but their approaches differ fundamentally from LTMs since they work on post-training of traditional autoregressive models on finetuning sets, while LTMs’ posterior inference is naturally optimized during pre-training. Sampling from $p_{\beta}({\mathbf{y}}|{\mathbf{x}},{\mathbf{z}})$ follows standard autoregressive sampling techniques (Freitag & Al-Onaizan, 2017; Holtzman et al., 2019). For unconditional generation, we sample from:
@@
An alternative sampling scheme is to incorporate each newly generated token into the prefix and then updating ${\mathbf{z}}$ through variational inference. We leave exploration of this more computationally intensive approach to future work.
@@
2.4 Inference-Time Computation
Compared to language models operating in the token space (e.g., ARMs and DDMs), LTMs introduce a distinct computational cost in the form of inference-time compute — a requirement stemming from the fast learning of latent thought vectors. This inference-time computation occurs in both model training and testing. Let’s start from analyzing it within the context of total training compute.
For one single iteration of LTM’s dual-rate learning with $T_{\text{fast}}$ inference steps on an input sequence of $N$ tokens (vocabulary size $V$), we consider a model with $L$ attention layers, $N_{{\mathbf{z}}}$ latent thought vectors per layer, and hidden dimension $H$. The forward pass computational complexity is approximately $\mathcal{O}(L(N^{2}H+NN_{{\mathbf{z}}}H+NH^{2})+NVH)$, comprising $\mathcal{O}(LN^{2}H)$ for self-attention, $\mathcal{O}(LNN_{{\mathbf{z}}}H)$ for cross-attention with latent vectors, $\mathcal{O}(LNH^{2})$ for feed-forward layers, and $\mathcal{O}(NVH)$ for embedding layers. The backward pass doubles this cost due to gradient computation and activation storage (Chowdhery et al., 2023). With $T_{\text{fast}}$ backward passes in fast learning, and $1_{\text{slow}}$ additional backward pass in slow learning, the training compute per token (trFLOPs/tok) is $\mathcal{O}((T_{\text{fast}}+1_{\text{slow}})L(N^{2}H+NN_{{\mathbf{z}}}H+NH^{2})+(T_{\text{fast}}+1_{\text{slow}})NVH)$. Thus, while both LTMs and ARMs involve gradient back-propagation for training, LTMs distribute compute differently: they trade ARMs’ compute in slow learning of global parameters for fast learning of local parameters.
[…]
During testing, $N$ varies by task: it represents the token sequence length for latent vector inference in likelihood estimation and generation tasks. As detailed in Section 2.3, generation tasks’ inference-time compute can further vary by sampling scheme. For our adopted sampling scheme, the trFLOPs/tok derived above provides a worst-case estimate of inference-time compute across all tasks.